OpenStudy (cmtboy2):

The formula gives the height h of a parallelogram with area A and base b. What is the height of a parallelogram with an area of 78 cm2 and a base of 13 cm?

OpenStudy (unklerhaukus):

Where is the formula?

OpenStudy (cmtboy2):

h=a/b

OpenStudy (unklerhaukus):

\(h=A/b\) substitute \(A\) = 78cm^2 and \(b\) = 12cm

OpenStudy (cmtboy2):

ok so add b and a

OpenStudy (unklerhaukus):

there is no addition sign, `/` is for division

OpenStudy (cmtboy2):

ok so i have to divid them right

OpenStudy (unklerhaukus):

\[h= \frac Ab=\frac{78\,\text{cm}^2}{12\,\text{cm}}=\frac{78}{12}\frac{\text{cm}^2}{\text{cm}}=\]

OpenStudy (unklerhaukus):

ops... 12 = 13

OpenStudy (unklerhaukus):

try with 13

OpenStudy (unklerhaukus):

78/13 = ? cm^2/cm = ??

OpenStudy (unklerhaukus):

do you know your 13 times tables?

OpenStudy (cmtboy2):

78/13=6

OpenStudy (cmtboy2):

yes i do

OpenStudy (unklerhaukus):

good, and what is cm^2 / cm

OpenStudy (cmtboy2):

6

OpenStudy (unklerhaukus):

so you have the number, but what is the units?

OpenStudy (cmtboy2):

?

OpenStudy (unklerhaukus):

cm^2 / cm = cm*cm/cm =

OpenStudy (cmtboy2):

6cm

OpenStudy (unklerhaukus):

yes

OpenStudy (cmtboy2):

one min i have more

OpenStudy (cmtboy2):

Solve the word problem. The height of a triangle with base b and area A is given by the formula h = . What is the height of a triangle with area 26 cm2 and base 8 cm? A. 1.625 cm B. 3.25 cm C. 3.5 cm D. 6.5 cm

OpenStudy (unklerhaukus):

For a triangle \[A = \tfrac12bh\] so \[h=2A/b\]

OpenStudy (unklerhaukus):

now A = 26 cm2 and b = 8 cm

OpenStudy (cmtboy2):

times a and b